### v1

# M4GT-Bench: Evaluation Benchmark for Black-Box Machine-Generated Text Detection Аннотация

M4GT-Bench: Evaluation Benchmark for Black-Box
Machine-Generated Text Detection
Аннотация
[…]
M4GT-Bench: Evaluation Benchmark for Black-Box
Machine-Generated Text Detection
Yuxia Wang,† Jonibek Mansurov,†, Petar Ivanov,† Jinyan Su,† Artem Shelmanov,† Akim Tsvigun,† Osama Mohammed Afzal,† Tarek Mahmoud,† Giovanni Puccetti,§ Thomas Arnold,¶ Alham Fikri Aji,† Nizar Habash,†‡ Iryna Gurevych,† Preslav Nakov† †Mohamed bin Zayed University of Artificial Intelligence, UAE ¶TU Darmstadt, Germany §Institute of Information Science and Technology, Italy ‡New York University Abu Dhabi, UAE {yuxia.wang, jonibek.mansurov, preslav.nakov}@mbzuai.ac.ae
utf8
M4GT-Bench: Evaluation Benchmark for Black-Box
Machine-Generated Text Detection
Yuxia Wang,† Jonibek Mansurov,†, Petar Ivanov,† Jinyan Su,† Artem Shelmanov,† Akim Tsvigun,† Osama Mohammed Afzal,† Tarek Mahmoud,† Giovanni Puccetti,§ Thomas Arnold,¶ Alham Fikri Aji,† Nizar Habash,†‡ Iryna Gurevych,† Preslav Nakov† †Mohamed bin Zayed University of Artificial Intelligence, UAE ¶TU Darmstadt, Germany §Institute of Information Science and Technology, Italy ‡New York University Abu Dhabi, UAE {yuxia.wang, jonibek.mansurov, preslav.nakov}@mbzuai.ac.ae
1 Introduction
The advent of Large Language Models (LLMs) such as ChatGPT 111We refer to GPT-3.5-Turbo throughout this paper. and GPT-4 marks a transformative era in text generation. These models are able to generate a coherent text that is very similar to human-written content. They are also easily accessible and becoming more and more widespread. While such tools significantly boost productivity across various fields such as journalism, social media, education, and academic writing, they also introduce unprecedented avenues for misuse and, consequently, pose negative societal implications. LLMs can generate immense amounts of deceptive fake news cluttering the information space. In social media, they have the potential to automatically generate fake accounts and increase their influence in social media communities without much effort from their owners. In the academic context, these technologies may advance beyond mere replication of existing content, potentially diminishing the effectiveness of established plagiarism detection methodologies and, consequently, posing a threat to the fundamental tenets of scholarly integrity. As a result, it could impact academic writing by leading to an increase in publications of automatically generated papers lacking scientific merit in respected journals and conferences. Finally, the scientific community has started reporting on cases of potentially automatically generated reviews in some well-known conferences. Studies have shown that humans perform only marginally better than random chance in distinguishing human-written from machine-generated texts (MGTs) Wang et al. (2023). Therefore, automatic MGT detection becomes essential in tackling misinformation, preserving the integrity of digital platforms and the scientific community, and ensuring trust in communications. Previous work typically frames MGT detection as binary classification Zellers et al. (2019); Mitchell et al. (2023); Bao et al. (2023), focusing primarily on English. Moreover, the majority of studies also overlook the fact that the content can be a mixture of an MGT and human-written text. These limitations make the previous experimental setups significantly different from practice. In this work, we address these issues by presenting a new benchmark for MGT detection that consists of three tasks, which were not previously explored in this exact formulation, each shedding light on different facets of this challenge.
Task 1: Binary Human-Written vs. Machine-Generated Text Classification. The objective is to categorize a given text as either human-written or machine-generated. This task is similar to the problem formulations in the previous work. However, our dataset offers annotated resources of new domains and generators in multiple languages. Task 1 sets two distinct tracks, one focusing exclusively on English sources (monolingual), and the other embracing a multilingual scope and introducing greater diversity compared to the previous work.
Task 2: Multi-Way Machine-Generated Text’s Generator Detection. This task challenges the identification of the specific language model responsible for generating a given full text. The text may be either human-written or crafted by a particular LLM. A different perspective to look at task 2 is that we are attributing authorship to a specific model Munir et al. (2021).
Task 3: Human-Written to Machine-Generated Text Change Point Detection. In this task, the goal is to precisely estimate the boundary within the mixed text, where the transition occurs from human-written to machine-generated content.
Our contributions are summarized below:
We introduce a novel task formulation for MGT detection, where human-machine mixed text is explored. It is the first attempt to evaluate the ability to detect a boundary between human-written and machine-generated texts.
We construct the most diverse public multilingual MGT detection benchmark involving nine languages, six domains, nine LLM generators and three different tasks.
Human evaluation utilizing few-shot learning, yields a performance level below random chance in distinguishing different LLM generations, suggesting the challenges of task 2. We further demonstrate strong baselines for each task and investigate their performance in various experimental settings.
2 Related Work
The task of detecting machine-generated text has long been formulated as a binary classification problem Zellers et al. (2019); Gehrmann et al. (2019a); Solaiman et al. (2019); Ippolito et al. (2019), where the goal is to distinguish texts generated by certain language models from those authored by humans. However, with the rapid advancement of a wide range of LLMs such as GPT-4, Bard, Cohere and Claude, they can be easily used to generate texts through simple API calls. There is a growing need for more fine-grained classification that can not only identify the nature of the texts (i.e., whether it is machine-generated or human written), but also its specific source (i.e., which LLM generates it?). This is also known as authorship attribution Uchendu et al. (2020); Venkatraman et al. (2023); Rivera Soto et al. (2024).
Binary Detection
In general, there are mainly two types of machine-generated text detector: supervised approaches Wang et al. (2023); Uchendu et al. (2021); Zellers et al. (2019); Zhong et al. (2020); Liu et al. (2022) and unsupervised methods based on white-box features such as likelihood and log-rank Solaiman et al. (2019); Ippolito et al. (2019); Mitchell et al. (2023); Su et al. (2023); He et al. (2023); Hans et al. (2024), and watermarking Kirchenbauer et al. (2023); Zhao et al. (2023b, a). We focus on supervised approaches. Wang et al. (2023) provide the evaluation results based on several supervised detectors, including RoBERTa Liu et al. (2019), XLM-R Conneau et al. (2019), logistic regression classifier with GLTR features Gehrmann et al. (2019b), stylistic features Li et al. (2014), andNELA Horne et al. (2019) features. Similar work using supervised methods can also be seen in Guo et al. (2023); Hu et al. (2023); Xiong et al. (2024). However, most of them only consider binary detection.
| Source | Human | Parallel Data | Total | New test | ||||||
| Domain | Total= | Upsample+ | Parallel | davinci-003 | ChatGPT | Cohere | Dolly-v2 | BLOOMz | Machine | GPT-4 |
| OUTFOX | 16,272 | 13,272 | 3,000 | 3,000 | 3,000 | 3,000 | 3,000 | 3,000 | 15,000 | 3,000 |
| Wikipedia | 14,333 | 11,997 | 2,336 | 3,000 | 2,995 | 2,336 | 2,702 | 2,999 | 14,032 | 3,000 |
| Wikihow | 15,999 | 13,000 | 2,999 | 3,000 | 5,557 | 3,000 | 3,000 | 3,000 | 17,557 | 3,000 |
| Reddit ELI5 | 16,000 | 13,000 | 3,000 | 3,000 | 3,000 | 3,000 | 3,000 | 2,999 | 14,999 | 3,000 |
| arXiv abstract | 15,998 | 13,000 | 2,998 | 3,000 | 3,000 | 3,000 | 3,000 | 3,000 | 15,000 | 3,000 |
| PeerRead | 2,847 | 0 | 2,847 | 2,340 | 2,340 | 2,342 | 2,344 | 2,334 | 11,700 | 2,334 |
| Total | 65,177 | 50,997 | 14,180 | 14,340 | 16,892 | 13,678 | 14,046 | 14,332 | 73,288 | 14,344 |
Multi-Class Detection
The fine-grained multi-class classification problem is closely related to authorship attribution. Munir et al. (2021) find that texts generated from LMs contain distinguishable signals that can be used to attribute the source of texts. Uchendu et al. (2020) investigate three authorship attribution problems: (1) whether two texts are generated by the same method, (2) whether a text is generated by machine or human, and (3) which language model generated the texts. Venkatraman et al. (2023) examine if the principle that humans prefer to spread information evenly during language production can help capture the unique signature of each LLM and human author while Rivera Soto et al. (2024) leverages representations of writing styles.
Authorship Obfuscation
goes beyond the binary/multi-class classification lens and considers an adversarial setting of co-authorship of human and machine Macko et al. (2024). It is known that machine-generated detection methods are susceptible to authorship obfuscation attacks such as paraphrasing, back-translation, and human-machine collaboration mixture Crothers et al. (2022); Krishna et al. (2023); Shi et al. (2023); Koike et al. (2023). Gao et al. (2024) introduces a dataset with a mixture of machine and human written texts using operations such as polishing, completing Xie et al. (2023), rewriting Shu et al. (2023), adding natural noise Wang et al. (2021), and adapting Gero et al. (2022). Kumarage et al. (2023) use stylometric signals to quantify changes in tweets and detect if and when AI starts to generate tweets. In this work, we focus on detecting and localizing the change point from human-written segments to machine-generated segments. We construct a dataset consisting of peer reviews and student essays based on ChatGPT, GPT-4, LLaMA-2 (7B, 13B and 70B). Models using semantic features show promising results when training and test sets are from the same domain and generators.
3 Datasets and Metrics
This section describes the datasets for each task and the details of their creation. The data statistics is shown in Table 1, 2 and 3.
| Source/ | Data | Language | Total | Parallel Data | |||||
| Domain | License | Human | Human | davinci-003 | ChatGPT | Jais-30B | LLaMA-2-70B | Total | |
| Arabic-Wikipedia | CC BY-SA-3.0 | Arabic | 1,209,042 | 3,000 | – | 3,000 | – | 6,000 | |
| True & Fake News | MIT License | Bulgarian | 94,000 | 3,000 | 3,000 | 3,000 | – | 9,000 | |
| Baike/Web QA | MIT license | Chinese | 113,313 | 3,000 | 3,000 | 3,000 | – | 9,000 | |
| id_newspapers_2018 | CC BY-NC-SA-4.0 | Indonesian | 499,164 | 3,000 | – | 3,000 | – | 6,000 | |
| RuATD | Apache 2.0 license | Russian | 75,291 | 3,000 | 3,000 | 3,000 | – | 9,000 | |
| Urdu-news | CC BY 4.0 | Urdu | 107,881 | 3,000 | – | 3,000 | – | 6,000 | |
| News | Apache 2.0 | Arabic | 1,000 | 1,000 | – | 1,000 | 100 | – | 2,100 |
| CHANGE-it News | CC BY-NC-SA 4.0 | Italian | 127,402 | 3,000 | – | – | – | 3,000 | 6,000 |
| News | CC BY-NC-SA-4.0 | German | 10,000 | 3,000 | – | 3,000 | – | – | 6,000 |
| Wikipedia | CC BY-SA-3.0 | German | 2,882,103 | 3,000 | – | 3,000 | – | – | 6,000 |
| Total | – | – | 5,119,196 | 28,000 | 9,000 | 25,000 | 100 | 3,000 | 65,100 |
3.1 Human vs. Machine
The task 1 aims to distinguish human-written from machine-generated text — a binary classification task. Based on the M4 dataset (Wang et al., 2023), given a domain, we up-sampled human text to match the total number of machine-generated text to avoid data imbalance. After upsampling, the dataset consists of 65,177 human-written texts and 73,288 machine-generated texts as shown in Table 1. We additionally generated texts using GPT-4 for each domain to evaluate the detectors’ generalization ability on unseen and strong generators. For multilingual track, we integrate new languages (German and Italian), and news for Arabic by generators of ChatGPT and Jais-30B, as the statistics shown in Table 2. Data Cleaning: simple artifacts in human-written texts, such as multiple newlines and bullet points, could assist detectors, as their presence in the training data may discourage detectors from learning more generalized signals. Therefore, we performed minimal cleaning of the upsampled human-written texts: (i) in a WikiHow text, we removed multiple commas at the beginning of a new line (like ‘‘,,,,,,,,,,, we believe that …’’) and repeating newlines (‘‘nnnnn text begin nnnnn”); (ii) in Wikipedia articles, we removed references (e.g., [1], [2]), URLs, multiple newlines, as well as paragraphs whose length was less than 50 characters; and (iii) in arXiv abstracts, we removed newlines stemming from PDF conversion. Metrics: in addition to accuracy, we also use precision, recall, and F1-score with respect to MGT.
3.2 Multi-way Detection
Task 2 is to determine which generator model produces the considered text. Different from task 1, we use the parallel subset for task 2 without the upsampling of human-written texts, including six generators: ChatGPT, davinci-003, GPT-4, Cohere, Dolly-v2, and BLOOMz. A new domain OUTFOX is collected to evaluate the classifier’s domain generalization in student essays. Metrics: F1-score, precision, recall in terms of seven individual labels representing different generators and human-written texts are used. We also provide F1-macro and accuracy.
| Domain | Generator | Train | Dev | Test | Total |
| PeerRead | ChatGPT | 3,649 (232) | 505 (23) | 1,522 (89) | 5,676 (344) |
| LLaMA-2-7B* | 3,649 (5) | 505 (0) | 1,035 (1) | 5,189 (6) | |
| LLaMA-2-7B | 3,649 (227) | 505 (24) | 1,522 (67) | 5,676 (318) | |
| LLaMA-2-13B | 3,649 (192) | 505 (24) | 1,522 (84) | 5,676 (300) | |
| LLaMA-2-70B | 3,649 (240) | 505 (21) | 1,522 (88) | 5,676 (349) | |
| OUTFOX | GPT-4 | – | – | 1,000 (10) | 1,000 (10) |
| LLaMA2-7B | – | – | 1,000 (8) | 1,000 (8) | |
| LLaMA2-13B | – | – | 1,000 (5) | 1,000 (5) | |
| LLaMA2-70B | – | – | 1,000 (19) | 1,000 (19) |
3.3 Boundary Identification
We consider two common misuse scenarios of human-machine mixed text — continuing to write academic paper reviews and student essays. Given the partial human-written essay with the corresponding problem statement and the partial human-written paper reviews with the title and abstract, we complete them by generating text based on GPT and LLaMA-2 series using prompts as shown in Figure 2. The proportion of words that are human-written range from 0 to 50%. The goal of a detection model is to identify the boundary position from which the text is generated by machines. Mixed texts are specifically generated using ChatGPT, GPT-4, and LLaMA-2 (7B, 13B and 70B). We generate 5,676 and 1,000 examples for two domains respectively using each generator (full statistics is presented in Table 3). Metrics: the Mean Absolute Error (MAE) is used to evaluate the performance of the boundary detection model. It measures the average absolute difference between the predicted position index and the actual changing point.
4 Human Evaluation
We investigate the complexity of differentiating the language model used for creating machine-generated texts for human readers. Considering that annotators may lose patience to learn and summarize the patterns of seven distinct generators, we control the number of unique models they could identify and simplify to four-class classification with five demonstration examples (five-shot) as practice to learn before the formal annotation.
4.1 Data for Human Evaluation
There are 35 unique domain-generator combinations — five domains seven generators (six LLMs plus human-written texts) for task 2 data.
With the concern of annotators’ workload, we sample 4 examples for each combination, in total of 140 examples, split into four groups as Figure 1.
Without consideration of out-of-domain classification, we follow the best-performing setup: see all, classify all.
If each annotator needs to check all domains, they need to learn 20 demonstration examples, which is a time consuming and laborious undertaking. Let us consider that each annotator annotates 3 domains. They would need to check 4 models 3 domains 4 examples = 48, which is in a more reasonable range.
Five demonstration examples are randomly selected, to ensure that four classes are all included.
Annotator background.
Group 1, 2, 3, and 4 are annotated by native Italian, Chinese, English, and Russian speakers, respectively; they are either postdoc or PhD students in NLP.
| Annotator | Prec | Recall | F1-macro | Accuracy |
| Group 1 | 27.91 | 28.25 | 28.18 | 27.42 |
| Group 2 | 16.28 | 09.22 | 12.91 | 10.45 |
| Group 3 | 20.93 | 16.00 | 16.91 | 15.47 |
| Group 4 | 25.58 | 25.83 | 25.25 | 24.82 |
| Group 1 + 2 | 22.09 | 19.73 | 21.10 | 20.27 |
| Group 3 + 4 | 23.26 | 22.98 | 20.84 | 21.13 |
| All | 22.67 | 22.11 | 21.06 | 21.20 |
4.2 Results
Intuitively, it would be challenging to distinguish between generators based on 1-shot per class. Annotators may make decisions depending on the observations that generations by Cohere tend to be repetitive, Dolly-v2 has high-quality and coherent texts, while BLOOMz texts are more likely to contain a bunch of numbers or a bunch of ‘‘YES’’ or ‘‘NO’’ characters repetitively at the end of the texts, compared to other models. For GPT series, annotators distinguish them depending on their quality level assumption that GPT-4 > ChatGPT > davinci-003. They spot human-written texts based on formatting patterns (e.g., initial double newlines, initial space, completely missing new lines in the paragraph), typos, inconsistencies within texts, or specific references and URLs. As shown in Table 4, the best distinction accuracy for humans is 27%, and the average accuracy 21% is less than random guess. This implies that it is extremely difficult for humans to learn patterns from five demonstration examples and then distinguish generations of different LLM generators (see the F1-score for each separate class in Appendix B).
5 Experiments
We applied two detectors for boundary identification based on semantic features. Five baseline classifiers suggested by Wang et al. (2023) for tasks 1 and 2 include: fine-tuned RoBERTa and XLM-R classifiers, logistic regression with GLTR features (LR-GLTR), SVM with Stylistic (Stylistic-SVM), and NELA features (NELA-SVM).
RoBERTa and XLM-R
We fine-tune two Transformer-based models for MGT detection: RoBERTa and XLM-R.
GLTR features
are based on the assumption that in order to generate fluent and natural-looking text, most LLM decoding strategies sample high-probability tokens from the head of the distribution Gehrmann et al. (2019a). Thus, word ranking information of an LM can be used to distinguish machine-generated texts from human-written. Following Wang et al. (2023), we select two categories of these features: (1) the number of tokens in the top-10, top-100, top-1000, and 1000+ ranks from the LM predicted probability distributions (4 features); and (2) the Frac(p) distribution over 10 bins ranging from 0.0 to 1.0 (10 features). Frac(p) describes the fraction of probability for the actual word divided by the maximum probability of any word at this position. We train a logistic regression model based on these 14 features to perform the binary classification task with the parameter of the maximum iteration = 1,000.
Stylistic and NELA features.
We use stylistic features including (1) character-based features, e.g., the number of characters, letters, special characters, etc., (2) syntactic features, e.g., a number of punctuation and function words, (3) structural features, e.g., a total number of sentences, and (4) word-based features, e.g., a total number of words, average word length, average sentence length, etc. NEws LAndscape (NELA) features (Horne et al., 2019) involves six aspects: (1) style: the style and the structure of the article; (2) complexity: how complex the writing is; (3) bias: overall bias and subjectivity; (4) affect: sentiment and emotional patterns; (5) moral based on the Moral Foundation Theory (Graham et al., 2012), and (6) event: time and location. An SVM classifier is applied to perform the binary distinction.
| Detector | Test | Prec | Recall | F1-score | Acc |
| RoBERTa | All | 99.16 | 99.56 | 99.36 | 99.26 |
| davinci-003 | 71.08 | 98.53 | 82.58 | 79.21 | |
| ChatGPT | 74.64 | 99.93 | 85.45 | 82.99 | |
| GPT-4 | 70.81 | 100.00 | 82.90 | 79.37 | |
| Cohere | 70.11 | 98.50 | 81.91 | 78.24 | |
| Dolly-v2 | 69.95 | 97.46 | 81.44 | 77.78 | |
| BLOOMz | 60.31 | 60.16 | 60.22 | 60.30 | |
| XLM-R | All | 95.08 | 98.80 | 96.87 | 96.31 |
| davinci-003 | 80.57 | 90.46 | 85.23 | 84.32 | |
| ChatGPT | 78.12 | 99.95 | 87.57 | 85.62 | |
| GPT-4 | 69.44 | 99.93 | 81.93 | 77.95 | |
| Cohere | 79.59 | 97.98 | 87.74 | 86.23 | |
| Dolly-v2 | 76.77 | 84.58 | 80.40 | 79.43 | |
| BLOOMz | 73.98 | 72.16 | 72.74 | 73.07 | |
| GLTR-LR | All | 84.59 | 88.71 | 86.60 | 84.26 |
| davinci-003 | 81.62 | 78.13 | 79.83 | 80.27 | |
| ChatGPT | 82.00 | 96.33 | 88.59 | 87.59 | |
| GPT-4 | 83.07 | 98.17 | 89.99 | 89.08 | |
| Cohere | 82.98 | 99.27 | 90.40 | 89.46 | |
| Dolly-v2 | 81.10 | 72.24 | 76.42 | 77.70 | |
| BLOOMz | 76.02 | 50.45 | 60.65 | 67.27 | |
| Stylistic-LR | All | 86.42 | 76.67 | 81.25 | 84.91 |
| davinci-003 | 81.14 | 47.16 | 59.65 | 68.10 | |
| ChatGPT | 65.96 | 50.67 | 57.31 | 62.26 | |
| GPT-4 | 97.62 | 44.82 | 61.44 | 71.87 | |
| Cohere | 75.67 | 44.18 | 55.79 | 64.99 | |
| Dolly-v2 | 77.75 | 49.37 | 60.39 | 67.62 | |
| BLOOMz | 57.93 | 48.86 | 53.01 | 56.69 | |
| NELA-LR | All | 74.55 | 63.78 | 68.75 | 75.27 |
| davinci-003 | 77.52 | 60.25 | 67.80 | 71.39 | |
| ChatGPT | 82.12 | 59.68 | 69.12 | 73.34 | |
| GPT-4 | 93.30 | 56.60 | 70.46 | 76.27 | |
| Cohere | 74.82 | 58.32 | 65.55 | 69.35 | |
| Dolly-v2 | 64.78 | 62.28 | 63.50 | 64.21 | |
| BLOOMz | 44.64 | 77.58 | 56.67 | 40.69 |
5.1 Monolingual Human vs. Machine
For task 1, we simulate a practical scenario where a detector have to deal with texts generated from a new LLM while it is trained on the outputs from a limited number of other generators. Practical detector should be able to generalize over different generators and be robust on unseen inputs.
Experiment Setup
We conduct experiments in two setups. (1) Given a single LLM, we combine its generations across all domains and the upsampled human-written texts as the test set. Texts produced by other LLMs and humans compile training and development sets (split by 4:1). (2) Over all data, we randomly split train and test data across all generators (row All in Table 5). We train detectors with five different seeds for all experiments and present the mean values of the metrics.
Results and Analysis
Classification results of five detectors are presented in Table 5. We compare results from three aspects: (1) given a detector, on which unseen generator, it generalizes the best (bold F1-score and Acc); (2) given a setting that training without generations of the generator to test, which detector performs the best (underlined Acc); and (3) on average across different settings, which detector is the most accurate and has the best generalization capabilities? Which detector has higher recall than precision? RoBERTa, XLM-R, and LR based on stylistic features give the best F1-score and accuracy on the setting where the classifier is trained with generations of all generators and tested across all generators as well. GLTR features perform best in the setting in which the detector is trained over data without generations from Cohere, and tested over Human vs. Cohere data, while NELA has the highest accuracy in distinguishing human vs. GPT-4 when it has not seen any data from GPT-4. The results for RoBERTa and XLM-R suggest that such detectors encountered notable challenges when confronted with text from new generators not encompassed in their training dataset, demonstrating the large accuracy gap between the setting of All and others. This problem is especially conspicuous in the case of BLOOMz for all detectors (All vs. BLOOMz), followed by Dolly-v2, davinci-003, ChatGPT, GPT-4 and Cohere. XLM-R and GLTR demonstrate better generalization performance (underlines numbers) over (davinci-003, Dolly-v2, and BLOOMz) and (ChatGPT, GPT-4, and Cohere) respectively. XLM-R is the best on average. It has higher accuracy and better generalization ability. LR with NELA features is the worst. Recall of RoBERTa, XLM-R, and GLTR is higher than precision, while the opposite for stylistic and NELA features. In the scenario of detecting misuse, we expect higher recall to avoid less misses of potentially malice.
5.2 Multilingual Human vs. Machine
Similar to monolingual setups, we evaluate detectors generalization on new languages (the language to test) and train on the combination of other languages. In the setting of All in Table 6, we merge data of all languages, and randomly split into train, validation and test sets.
Results and Analysis
We categorize the performance of the detector (XLM-R) across ten configurations into three tiers based on the F1 score. (1) F1<50: Bulgarian, Russian and Indonesian; (2) 60<F1<85: English, German, Italian, Chinese; and (3) F1>90: Arabic, Urdu, All. For the first level, Bulgarian and Russian are both Slavic languages, and Indonesian is a member of the Malayo-Polynesian branch of the language family, training on other distantly-unrelated languages (Latin, Chinese and Arabic family) would naturally result in low performance. For the second level, Chinese is a high-resource language, so the detector could benefit from the learned patterns during XLM-R pretraining. For German and Italian, English training data will keep the same family languages in a medium performance, vice versa. High recall and low precision with respect to MGT across all testing languages implies that the majority of examples are recognized as machine-generated text. This leads to lower accuracy on human-written text than that on MGT (unweighted average macro-F1 < MGT F1). Overall, the detector often predicts texts as MGT. when dealing with low-resource languages or training on language families that are distantly related, the accuracy tends to be moderate.
| Test | Prec | Recall | F1-score | Acc | F1-macro |
| All | 91.49 | 98.52 | 94.86 | 94.52 | 94.49 |
| Arabic | 88.91 | 97.35 | 92.64 | 92.18 | 92.12 |
| Bulgarian | 51.41 | 99.97 | 67.90 | 52.74 | 39.18 |
| Chinese | 76.85 | 95.18 | 84.73 | 82.51 | 81.93 |
| English | 63.72 | 84.64 | 72.42 | 66.51 | 64.55 |
| German | 66.21 | 98.30 | 79.00 | 73.62 | 71.58 |
| Indonesian | 53.13 | 100.0 | 69.39 | 55.83 | 45.00 |
| Italian | 75.20 | 100.0 | 85.84 | 83.51 | 83.04 |
| Russian | 52.55 | 86.38 | 65.26 | 53.70 | 47.60 |
| Urdu | 91.29 | 97.93 | 94.49 | 94.39 | 94.39 |
| Detector | Test | Overall Performance | |||
| Domain | Prec | Recall | F1-macro | Acc | |
| RoBERTa | All | 96.35 | 96.21 | 96.23 | 96.25 |
| arXiv | 60.55 | 41.11 | 33.67 | 41.11 | |
| PeerRead | 67.54 | 65.47 | 61.55 | 64.77 | |
| 73.59 | 67.97 | 62.33 | 67.97 | ||
| WikiHow | 70.94 | 67.60 | 64.91 | 64.20 | |
| Wikipedia | 40.44 | 50.10 | 37.61 | 49.64 | |
| XLM-R | All | 87.69 | 86.26 | 86.03 | 86.36 |
| arXiv | 51.40 | 41.73 | 38.82 | 41.73 | |
| PeerRead | 52.20 | 51.66 | 44.84 | 50.38 | |
| 69.82 | 60.58 | 59.92 | 60.58 | ||
| WikiHow | 64.01 | 60.81 | 57.38 | 56.13 | |
| Wikipedia | 58.57 | 41.07 | 37.92 | 40.25 | |
| GLTR-LR | All | 42.9 | 46.11 | 43.07 | 45.84 |
| arXiv | 21.5 | 30.61 | 23.19 | 30.60 | |
| PeerRead | 43.50 | 46.24 | 42.52 | 46.37 | |
| 46.88 | 47.91 | 43.90 | 47.91 | ||
| WikiHow | 43.87 | 38.78 | 38.28 | 37.23 | |
| Wikipedia | 41.91 | 35.66 | 35.54 | 34.08 | |
| GLTR-SVM | All | 53.19 | 53.47 | 52.50 | 53.75 |
| arXiv | 27.88 | 35.40 | 29.25 | 35.40 | |
| PeerRead | 33.98 | 37.18 | 34.27 | 37.54 | |
| 46.68 | 47.84 | 44.73 | 47.84 | ||
| WikiHow | 43.88 | 42.36 | 40.52 | 37.84 | |
| Wikipedia | 43.62 | 37.96 | 35.69 | 37.18 | |
| Stylistic-SVM | All | 78.57 | 37.80 | 47.86 | 35.53 |
| arXiv | 46.57 | 7.67 | 11.55 | 7.44 | |
| PeerRead | 53.40 | 27.16 | 30.39 | 25.30 | |
| 61.96 | 28.23 | 33.07 | 26.53 | ||
| WikiHow | 54.48 | 34.89 | 39.94 | 26.49 | |
| Wikipedia | 51.40 | 23.56 | 29.54 | 17.78 | |
| NELA-SVM | All | 66.99 | 27.48 | 34.80 | 25.60 |
| arXiv | 53.46 | 12.97 | 19.11 | 11.71 | |
| PeerRead | 48.53 | 24.76 | 22.84 | 23.31 | |
| 45.94 | 28.09 | 31.45 | 24.20 | ||
| WikiHow | 46.87 | 24.81 | 28.54 | 21.16 | |
| Wikipedia | 47.81 | 22.62 | 27.36 | 19.79 | |
| Detector | Train Data | Peerread LLaMA-2-7B* | Peerread ChatGPT | All Test |
| Longformer | All | 1.89 0.79 | 4.36 0.36 | 21.54 0.25 |
| ChatGPT | 31.43 6.15 | 4.55 0.36 | 25.14 0.93 | |
| LLaMA-2-7B* | 1.94 0.07 | 51.379 0.72 | 53.62 1.60 | |
| Deberta-v3 | All | 0.57 0.23 | 2.63 0.20 | 15.55 2.60 |
| ChatGPT | 14.96 2.19 | 2.53 0.09 | 19.67 1.05 | |
| LLaMA-2-7B* | 0.66 0.12 | 24.59 4.07 | 32.35 0.78 |
5.3 Multi-way Detection
Experimental Setup
For task 2, the classifiers are trained to distinguish multiple generators (seven-label classification), including human, davinci-003, ChatGPT, GPT-4, Cohere, Dolly-v2, BLOOMz. We measure the capabilities of detectors to generalize on a new domain that was not presented during training. The detectors are trained on the whole dataset except the domain used for testing. For this task, we calculate accuracy, macro F1-score, and the class-wise F1-scores. We experiment with six baselines: fine-tuned RoBERTa and XLM-R, logistic regression (one-vs-rest) and SVM (one-vs-one) with GLTR features, and SVM (one-vs-one) with stylistic and NELA.
Results and Analysis
We try to answer three questions: (1) given a detector, which domain is the hardest to predict; (2) which generator is the most challenging one to distinguish; and (3) which detector performs the best on task 2. As results in Table 7, we see that Transformer-based black-box detectors demonstrate poor performance in this task, when facing with text from new domains not included in their training data. The detectors exhibit significant challenges in generalizing to unfamiliar content, reflecting a notable weakness in their ability to handle previously unseen domains. The domain of arXiv is the hardest for most detectors to identify if they are trained without arXiv examples. For most domains, davinci-003 is the most challenging generator to distinguish followed by ChatGPT. BLOOMz appears to be the easiest one, for all classifiers (see class-wise F1-scores in Appendix C). Even the worst detector NELA-SVM can identify BLOOMz with >90% F1-score. This may result from the fact that the generative distribution of BLOOMz is significantly different from other generators, while the distribution between davinci-003 and ChatGPT are more similar to each other, making it hard to make distinctions. RoBERTa is the best in this task followed by XLM-R. Based on the GLTR features, SVM one-vs-one mode is better than logistic regression one-vs-rest mode. Under the same SVM, GLTR features are more helpful than stylistic and NELA.
5.4 Boundary Identification
Experimental Setup
We perform evaluation for task 3 in two settings: (1) train on PeerRead and test on PeerRead (same or different generators); and (2) train on PeerRead and test on a new domain with multiple generators (test column in Table 3 with 11,123 examples). As a baseline, we use a sequence tagger based on Longformer (Beltagy et al., 2020) considering the long context (> 1,024 tokens) and DeBERTa-v3 (He et al., 2021), where the position of the first word that is predicted as machine-generated is our prediction. The label was predicted at the token level: human-written words are tagged as 0 and machine-generated words are tagged as 1. We remap the tagged tokens back to the word level and calculate the MAE. We run experiments three times with different random seeds and calculate the mean and standard deviation.
Results and Analysis
The results for task 3 are presented in Table 8. When training and testing within the same domain, it is much harder for boundary detection models to predict changing points for unseen generators, e.g., when training on generations from LLaMA-2 and testing on ChatGPT and vice versa. Testing on an unseen domain (mixture of generators) results in worse results, from MAE<5 (under the setting of All for both detectors) to MAE>21 and > 15 for Longformer and Deterta-v3 respectively, especially for Longformer trained with LLaMA-2 data (MAE>53). This demonstrates challenges on unseen domains and generators, consistent with findings in task 1 and 2.
6 Conclusion and Future Direction
In this work, we extend M4 into a new benchmark M4GT-Bench for MGT detection, providing the widest coverage of nine languages, six domains, and nine state-of-the-art generators, including GPT-4 and LLaMA-2 series. We formulate three tasks, with a novel formulation of detecting the changing point from human-written to machine-generated. We observe that human readers are unable to consistently differentiate between generators and operate merely at the level of making random guesses. For all tasks, we find that the detector usually suffers a severe penalty in their performance when they are challenged with unseen domains or generators. Some generators such as BLOOMz and domains such as arXiv appear harder than others. Overall, we see that Transformer-based detectors such as RoBERTa and XLM-R are usually the best on average. In future work, we plan to develop a demonstration system assisting human to detect MGT, especially in soem high-stake domains such as clinical, finantial and laws, and distinguish the authorship of different LLMs, promoting the protection of intelligence property. In addition to text, we would extend to other modalities detection including speech, image and videos. From the perspective of adversarial attack such as paraphrasing, it is worthwhile to explore the mixed texts with multiple changing points or vague boundaries.
Limitations
While we provide strong baselines and analyze their performance in various evaluation settings, certain limitations remain for future work. Firstly, the nature of supervised multi-way classification of different LLMs disables the distinction of unseen generators, and tends to fall into random guess once the distribution of training and test data differs. Also, it is vulnerable to language style attack such as paraphrasing in different tones, back-translation and other textual adversarial attacks. Methods based on watermark and few-shot in-content learning are more promising for this task. Secondly, for all tasks, current black-box approaches may be less effective and may demonstrate poor generalization for unseen domains, generators, and languages, and this suggests the need to study more general methods to improve the detection and the potential misuse of LLMs. Thirdly, we simplify the task of boundary identification by assuming that the presented text is a mixture of first human-written segments and then machine-generated segments with only one changing point. Practical scenarios are more complicated. We should first detect whether it is a mixed text, and further identify all changing points.
Ethics and Broader Impact
We discuss some potential ethical concerns about the present work.
Data Collection and Licenses
We used pre-existing corpora: M4 and OUTFOX that have been publicly released and approved for research purposes, with clear dataset licenses. Data licenses for multilingual dataset are listed in Table 2.
Security Implication
The M4GT-Bench is intended for the development of a robust MGT detection system to identify and mitigate misuse, such as blocking the spread of automated misinformation campaigns, safeguarding individuals and institutions from potential financial losses. For domains of journalism, academia, and legal processes, the authenticity of information is paramount MGT detection ensures the integrity of content in these fields, thereby preserving trust. Also, users become more aware of LLMs capabilities, a growing skepticism towards digital content happens. Effective MGT detection can alleviate these concerns, ensuring users can trust LLM generations.
Список литературы
- Bao et al. (2023) Guangsheng Bao, Yanbin Zhao, Zhiyang Teng, Linyi Yang, and Yue Zhang. 2023. Fast-detectgpt: Efficient zero-shot detection of machine-generated text via conditional probability curvature. arXiv preprint arXiv:2310.05130.
- Beltagy et al. (2020) Iz Beltagy, Matthew E. Peters, and Arman Cohan. 2020. Longformer: The long-document transformer. CoRR, abs/2004.05150.
- Conneau et al. (2019) Alexis Conneau, Kartikay Khandelwal, Naman Goyal, Vishrav Chaudhary, Guillaume Wenzek, Francisco Guzmán, Edouard Grave, Myle Ott, Luke Zettlemoyer, and Veselin Stoyanov. 2019. Unsupervised cross-lingual representation learning at scale. arXiv preprint arXiv:1911.02116.
- Crothers et al. (2022) Evan Crothers, Nathalie Japkowicz, Herna Viktor, and Paula Branco. 2022. Adversarial robustness of neural-statistical features in detection of generative transformers. In 2022 International Joint Conference on Neural Networks (IJCNN), pages 1–8. IEEE.
- Gao et al. (2024) Chujie Gao, Dongping Chen, Qihui Zhang, Yue Huang, Yao Wan, and Lichao Sun. 2024. Llm-as-a-coauthor: The challenges of detecting llm-human mixcase. arXiv preprint arXiv:2401.05952.
- Gehrmann et al. (2019a) Sebastian Gehrmann, Hendrik Strobelt, and Alexander Rush. 2019a. GLTR: Statistical detection and visualization of generated text. In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics: System Demonstrations, pages 111–116, Florence, Italy. Association for Computational Linguistics.
- Gehrmann et al. (2019b) Sebastian Gehrmann, Hendrik Strobelt, and Alexander M Rush. 2019b. Gltr: Statistical detection and visualization of generated text. arXiv preprint arXiv:1906.04043.
- Gero et al. (2022) Katy Ilonka Gero, Vivian Liu, and Lydia Chilton. 2022. Sparks: Inspiration for science writing using language models. In Designing interactive systems conference, pages 1002–1019.
- Graham et al. (2012) Jesse Graham, Jonathan Haidt, Sena Koleva, Matt Motyl, Ravi Iyer, Sean Wojcik, and Peter Ditto. 2012. Moral foundations theory: The pragmatic validity of moral pluralism. Advances in Experimental Social Psychology, 47.
- Guo et al. (2023) Biyang Guo, Xin Zhang, Ziyuan Wang, Minqi Jiang, Jinran Nie, Yuxuan Ding, Jianwei Yue, and Yupeng Wu. 2023. How close is chatgpt to human experts? comparison corpus, evaluation, and detection. CoRR, abs/2301.07597.
- Hans et al. (2024) Abhimanyu Hans, Avi Schwarzschild, Valeriia Cherepanova, Hamid Kazemi, Aniruddha Saha, Micah Goldblum, Jonas Geiping, and Tom Goldstein. 2024. Spotting llms with binoculars: Zero-shot detection of machine-generated text. arXiv preprint arXiv:2401.12070.
- He et al. (2021) Pengcheng He, Xiaodong Liu, Jianfeng Gao, and Weizhu Chen. 2021. Deberta: decoding-enhanced bert with disentangled attention. In 9th International Conference on Learning Representations, ICLR 2021, Virtual Event, Austria, May 3-7, 2021. OpenReview.net.
- He et al. (2023) Xinlei He, Xinyue Shen, Zeyuan Chen, Michael Backes, and Yang Zhang. 2023. Mgtbench: Benchmarking machine-generated text detection. CoRR, abs/2303.14822.
- Horne et al. (2019) Benjamin D Horne, Jeppe Nørregaard, and Sibel Adali. 2019. Robust fake news detection over time and attack. ACM Transactions on Intelligent Systems and Technology (TIST), 11(1):1–23.
- Hu et al. (2023) Xiaomeng Hu, Pin-Yu Chen, and Tsung-Yi Ho. 2023. Radar: Robust ai-text detection via adversarial learning. arXiv preprint arXiv:2307.03838.
- Ippolito et al. (2019) Daphne Ippolito, Daniel Duckworth, Chris Callison-Burch, and Douglas Eck. 2019. Automatic detection of generated text is easiest when humans are fooled. arXiv preprint arXiv:1911.00650.
- Kirchenbauer et al. (2023) John Kirchenbauer, Jonas Geiping, Yuxin Wen, Jonathan Katz, Ian Miers, and Tom Goldstein. 2023. A watermark for large language models. CoRR, abs/2301.10226.
- Koike et al. (2023) Ryuto Koike, Masahiro Kaneko, and Naoaki Okazaki. 2023. Outfox: Llm-generated essay detection through in-context learning with adversarially generated examples. arXiv preprint arXiv:2307.11729.
- Krishna et al. (2023) Kalpesh Krishna, Yixiao Song, Marzena Karpinska, John Wieting, and Mohit Iyyer. 2023. Paraphrasing evades detectors of ai-generated text, but retrieval is an effective defense. arXiv preprint arXiv:2303.13408.
- Kumarage et al. (2023) Tharindu Kumarage, Joshua Garland, Amrita Bhattacharjee, Kirill Trapeznikov, Scott Ruston, and Huan Liu. 2023. Stylometric detection of ai-generated text in twitter timelines. arXiv preprint arXiv:2303.03697.
- Li et al. (2014) Jenny S Li, John V Monaco, Li-Chiou Chen, and Charles C Tappert. 2014. Authorship authentication using short messages from social networking sites. In 2014 IEEE 11th International Conference on e-Business Engineering, pages 314–319. IEEE.
- Liu et al. (2022) Xiaoming Liu, Zhaohan Zhang, Yichen Wang, Hang Pu, Yu Lan, and Chao Shen. 2022. Coco: Coherence-enhanced machine-generated text detection under data limitation with contrastive learning. arXiv preprint arXiv:2212.10341.
- Liu et al. (2019) Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov. 2019. Roberta: A robustly optimized bert pretraining approach. arXiv preprint arXiv:1907.11692.
- Macko et al. (2024) Dominik Macko, Robert Moro, Adaku Uchendu, Ivan Srba, Jason Samuel Lucas, Michiharu Yamashita, Nafis Irtiza Tripto, Dongwon Lee, Jakub Simko, and Maria Bielikova. 2024. Authorship obfuscation in multilingual machine-generated text detection. arXiv preprint arXiv:2401.07867.
- Mitchell et al. (2023) Eric Mitchell, Yoonho Lee, Alexander Khazatsky, Christopher D. Manning, and Chelsea Finn. 2023. Detectgpt: Zero-shot machine-generated text detection using probability curvature. CoRR, abs/2301.11305.
- Munir et al. (2021) Shaoor Munir, Brishna Batool, Zubair Shafiq, Padmini Srinivasan, and Fareed Zaffar. 2021. Through the looking glass: Learning to attribute synthetic text generated by language models. In Proceedings of the 16th Conference of the European Chapter of the Association for Computational Linguistics: Main Volume, pages 1811–1822.
- Rivera Soto et al. (2024) Rafael Rivera Soto, Kailin Koch, Aleem Khan, Barry Chen, Marcus Bishop, and Nicholas Andrews. 2024. Few-shot detection of machine-generated text using style representations. arXiv e-prints, pages arXiv–2401.
- Shi et al. (2023) Zhouxing Shi, Yihan Wang, Fan Yin, Xiangning Chen, Kai-Wei Chang, and Cho-Jui Hsieh. 2023. Red teaming language model detectors with language models. arXiv preprint arXiv:2305.19713.
- Shu et al. (2023) Lei Shu, Liangchen Luo, Jayakumar Hoskere, Yun Zhu, Canoee Liu, Simon Tong, Jindong Chen, and Lei Meng. 2023. Rewritelm: An instruction-tuned large language model for text rewriting. arXiv preprint arXiv:2305.15685.
- Solaiman et al. (2019) Irene Solaiman, Miles Brundage, Jack Clark, Amanda Askell, Ariel Herbert-Voss, Jeff Wu, Alec Radford, Gretchen Krueger, Jong Wook Kim, Sarah Kreps, et al. 2019. Release strategies and the social impacts of language models. arXiv preprint arXiv:1908.09203.
- Su et al. (2023) Jinyan Su, Terry Yue Zhuo, Di Wang, and Preslav Nakov. 2023. Detectllm: Leveraging log rank information for zero-shot detection of machine-generated text. arXiv preprint arXiv:2306.05540.
- Uchendu et al. (2020) Adaku Uchendu, Thai Le, Kai Shu, and Dongwon Lee. 2020. Authorship attribution for neural text generation. In Proceedings of the 2020 conference on empirical methods in natural language processing (EMNLP), pages 8384–8395.
- Uchendu et al. (2021) Adaku Uchendu, Zeyu Ma, Thai Le, Rui Zhang, and Dongwon Lee. 2021. Turingbench: A benchmark environment for turing test in the age of neural text generation. arXiv preprint arXiv:2109.13296.
- Venkatraman et al. (2023) Saranya Venkatraman, Adaku Uchendu, and Dongwon Lee. 2023. Gpt-who: An information density-based machine-generated text detector. arXiv preprint arXiv:2310.06202.
- Wang et al. (2021) Boxin Wang, Chejian Xu, Shuohang Wang, Zhe Gan, Yu Cheng, Jianfeng Gao, Ahmed Hassan Awadallah, and Bo Li. 2021. Adversarial glue: A multi-task benchmark for robustness evaluation of language models. arXiv preprint arXiv:2111.02840.
- Wang et al. (2023) Yuxia Wang, Jonibek Mansurov, Petar Ivanov, Jinyan Su, Artem Shelmanov, Akim Tsvigun, Chenxi Whitehouse, Osama Mohammed Afzal, Tarek Mahmoud, Alham Fikri Aji, et al. 2023. M4: Multi-generator, multi-domain, and multi-lingual black-box machine-generated text detection. arXiv preprint arXiv:2305.14902.
- Xie et al. (2023) Zhuohan Xie, Trevor Cohn, and Jey Han Lau. 2023. The next chapter: A study of large language models in storytelling. In Proceedings of the 16th International Natural Language Generation Conference, pages 323–351.
- Xiong et al. (2024) Feng Xiong, Thanet Markchom, Ziwei Zheng, Subin Jung, Varun Ojha, and Huizhi Liang. 2024. Fine-tuning large language models for multigenerator, multidomain, and multilingual machine-generated text detection. arXiv preprint arXiv:2401.12326.
- Zellers et al. (2019) Rowan Zellers, Ari Holtzman, Hannah Rashkin, Yonatan Bisk, Ali Farhadi, Franziska Roesner, and Yejin Choi. 2019. Defending against neural fake news. In Advances in Neural Information Processing Systems 32: Annual Conference on Neural Information Processing Systems 2019, NeurIPS 2019, December 8-14, 2019, Vancouver, BC, Canada, pages 9051–9062.
- Zhao et al. (2023a) Xuandong Zhao, Prabhanjan Ananth, Lei Li, and Yu-Xiang Wang. 2023a. Provable robust watermarking for ai-generated text. arXiv preprint arXiv:2306.17439.
- Zhao et al. (2023b) Xuandong Zhao, Yu-Xiang Wang, and Lei Li. 2023b. Protecting language generation models via invisible watermarking. CoRR, abs/2302.03162.
- Zhong et al. (2020) Wanjun Zhong, Duyu Tang, Zenan Xu, Ruize Wang, Nan Duan, Ming Zhou, Jiahai Wang, and Jian Yin. 2020. Neural deepfake detection with factual structure of text. arXiv preprint arXiv:2010.07475.
Appendix
Приложение A Task 3 Template Prompt
Figure 2 provides template prompts used for task 3, instructing LLMs to continue to write peer reviews and student essays with human-written part as the context.
Приложение B Human Evaluation Results
Table 9 presents the overall performance and generator-wise accuracy for human evaluation on task 2.
| Classifier | Overall Performance | Separate class F1-score | |||||||||
| Prec | Recall | F1-macro | Acc | Human | davinci-003 | ChatGPT | GPT-4 | Cohere | Dolly-v2 | BLOOMz | |
| Group 1 | 27.91 | 28.25 | 28.18 | 27.42 | 12.50 | 32.00 | 23.08 | 42.11 | — | — | — |
| Group 2 | 16.28 | 09.22 | 12.91 | 10.45 | 32.26 | — | — | — | 10.00 | 10.00 | 0.00 |
| Group 3 | 20.93 | 16.00 | 16.91 | 15.47 | 32.26 | — | — | — | 9.52 | 13.33 | 22.22 |
| Group 4 | 25.58 | 25.83 | 25.25 | 24.82 | 25.00 | 47.62 | 0.00 | 26.67 | — | — | — |
| Group 1 + 2 | 22.09 | 19.73 | 21.10 | 20.27 | 25.53 | 32.00 | 22.22 | 42.11 | 10.00 | 10.00 | 0.00 |
| Group 3 + 4 | 23.26 | 22.98 | 20.84 | 21.13 | 28.57 | 47.62 | 0.00 | 26.67 | 9.52 | 13.33 | 22.22 |
| All | 22.67 | 22.11 | 21.06 | 21.20 | 27.27 | 39.13 | 13.04 | 35.29 | 9.76 | 11.43 | 12.50 |
Приложение C Results of Multi-way Classification
Table 10 provides the overall performance and class-wise (generator) F1-scores.
| Detector | Test | Overall Performance | Separate class F1-score | |||||||||
| Domain | Prec | Recall | F1-macro | Acc | Human | davinci-003 | ChatGPT | GPT-4 | Cohere | Dolly-v2 | BLOOMz | |
| RoBERTa | All | 96.35 | 96.21 | 96.23 | 96.25 | 96.34 | 94.43 | 96.38 | 96.74 | 97.12 | 93.30 | 99.31 |
| arXiv | 60.55 | 41.11 | 33.67 | 41.11 | 44.91 | 0.49 | 27.97 | 64.61 | 2.28 | 38.91 | 56.49 | |
| PeerRead | 67.54 | 65.47 | 61.55 | 64.77 | 58.71 | 0.47 | 66.45 | 53.79 | 62.60 | 94.25 | 94.60 | |
| 73.59 | 67.97 | 62.33 | 67.97 | 89.89 | 10.38 | 50.14 | 29.12 | 84.72 | 77.12 | 94.92 | ||
| WikiHow | 70.94 | 67.60 | 64.91 | 64.20 | 27.95 | 39.25 | 50.45 | 86.94 | 67.46 | 83.72 | 98.57 | |
| Wikipedia | 40.44 | 50.10 | 37.61 | 49.64 | 0.22 | 1.02 | 0.53 | 64.30 | 64.67 | 55.78 | 76.74 | |
| XLM-R | All | 87.69 | 86.26 | 86.03 | 86.36 | 91.41 | 80.63 | 87.48 | 90.30 | 81.60 | 72.03 | 98.75 |
| arXiv | 51.40 | 41.73 | 38.82 | 41.73 | 52.99 | 10.15 | 73.11 | 7.06 | 39.07 | 40.30 | 49.08 | |
| PeerRead | 52.20 | 51.66 | 44.84 | 50.38 | 17.19 | 0.53 | 66.21 | 32.68 | 38.51 | 62.64 | 96.13 | |
| 69.82 | 60.58 | 59.92 | 60.58 | 75.21 | 48.82 | 49.27 | 37.89 | 58.66 | 52.55 | 97.02 | ||
| WikiHow | 64.01 | 60.81 | 57.38 | 56.13 | 43.67 | 55.08 | 26.79 | 64.95 | 66.35 | 46.13 | 98.67 | |
| Wikipedia | 58.57 | 41.07 | 37.92 | 40.25 | 8.42 | 40.51 | 31.51 | 36.79 | 38.28 | 48.21 | 61.70 | |
| GLTR-LR | All | 42.9 | 46.11 | 43.07 | 45.84 | 66.98 | 13.15 | 37.06 | 30.17 | 50.30 | 33.83 | 70.00 |
| arXiv | 21.5 | 30.61 | 23.19 | 30.60 | 55.28 | 0.31 | 1.45 | 3.24 | 30.07 | 32.22 | 39.80 | |
| PeerRead | 43.50 | 46.24 | 42.52 | 46.37 | 58.77 | 1.05 | 36.18 | 50.92 | 43.31 | 39.51 | 67.90 | |
| 46.88 | 47.91 | 43.90 | 47.91 | 84.05 | 22.77 | 19.58 | 35.48 | 62.80 | 13.20 | 69.43 | ||
| WikiHow | 43.87 | 38.78 | 38.28 | 37.23 | 57.49 | 14.69 | 28.41 | 11.35 | 54.79 | 18.79 | 82.44 | |
| Wikipedia | 41.91 | 35.66 | 35.54 | 34.08 | 67.54 | 11.78 | 14.10 | 20.12 | 32.22 | 33.90 | 69.14 | |
| GLTR-SVM | All | 53.19 | 53.47 | 52.50 | 53.75 | 70.15 | 32.16 | 53.37 | 42.57 | 53.57 | 41.41 | 74.28 |
| arXiv | 27.88 | 35.40 | 29.25 | 35.40 | 61.14 | 0.06 | 42.95 | 0.00 | 28.55 | 33.7 | 38.33 | |
| PeerRead | 33.98 | 37.18 | 34.27 | 37.54 | 60.61 | 0.11 | 31.57 | 1.29 | 31.45 | 41.64 | 73.21 | |
| 46.68 | 47.84 | 44.73 | 47.84 | 83.34 | 14.95 | 43.89 | 18.88 | 56.88 | 15.58 | 79.62 | ||
| WikiHow | 43.88 | 42.36 | 40.52 | 37.84 | 66.13 | 22.54 | 1.15 | 23.16 | 64.09 | 23.65 | 82.93 | |
| Wikipedia | 43.62 | 37.96 | 35.69 | 37.18 | 53.55 | 17.18 | 40.28 | 43.48 | 32.03 | 36.89 | 26.45 | |
| Stylistic-SVM | All | 78.57 | 37.80 | 47.86 | 35.53 | 45.16 | 18.85 | 46.54 | 56.60 | 48.02 | 26.48 | 93.36 |
| arXiv | 46.57 | 7.67 | 11.55 | 7.44 | 6.39 | 0.00 | 2.03 | 2.99 | 3.45 | 9.87 | 56.10 | |
| PeerRead | 53.40 | 27.16 | 30.39 | 25.30 | 17.27 | 0.58 | 6.15 | 60.72 | 5.64 | 33.37 | 89.01 | |
| 61.96 | 28.23 | 33.07 | 26.53 | 50.18 | 15.22 | 10.78 | 26.11 | 29.06 | 12.92 | 87.18 | ||
| WikiHow | 54.48 | 34.89 | 39.94 | 26.49 | 43.20 | 26.8 | 29.77 | 38.6 | 25.26 | 23.34 | 92.60 | |
| Wikipedia | 51.40 | 23.56 | 29.54 | 17.78 | 33.78 | 8.07 | 1.01 | 38.78 | 33.45 | 34.5 | 57.18 | |
| NELA-SVM | All | 66.99 | 27.48 | 34.80 | 25.60 | 37.64 | 0.36 | 18.63 | 30.52 | 48.00 | 14.75 | 93.72 |
| arXiv | 53.46 | 12.97 | 19.11 | 11.71 | 13.64 | 0.07 | 12.88 | 10.73 | 19.00 | 10.81 | 66.62 | |
| PeerRead | 48.53 | 24.76 | 22.84 | 23.31 | 7.80 | 0.00 | 5.1 | 41.17 | 9.59 | 4.72 | 91.50 | |
| 45.94 | 28.09 | 31.45 | 24.20 | 49.00 | 0.46 | 8.69 | 8.32 | 48.28 | 13.96 | 91.43 | ||
| WikiHow | 46.87 | 24.81 | 28.54 | 21.16 | 48.68 | 0.20 | 6.96 | 13.88 | 16.80 | 16.65 | 96.59 | |
| Wikipedia | 47.81 | 22.62 | 27.36 | 19.79 | 31.19 | 0.07 | 6.88 | 5.79 | 29.46 | 24.95 | 93.19 | |
Приложение D Examples of Task 1 and 2
For English, we exhibit two examples from GPT-4-arXiv and GPT-4-OUTFOX. For Arabic, we present two examples generated by ChatGPT and Jais-30B respectively based on news. Afterwards, two examples generated by ChatGPT for German news and Wikipedia are demonstrated, followed by an Italian example produced by a fine-tuned LLaMA-2-70B using Italian corpus, the model named camoscio-70B.
| Field | Content |
| Language | English |
| LLM Generator | GPT-4 |
| Domain | OUTFOX |
| Problem Statement |
Explain the reasons why distractions while driving, specifically texting and driving, should be illegal and the consequences of not following this law. Use statistics and personal opinion to support your argument. |
| Human |
Distractions while driving could lead to deaths of other civilians. There are reasons why this topic became an actual law. The law was just passed last year that is you’re on your phone while driving, that is an immediate ticket. Due to your criminal record (if it’s bad) you can even do jail time if you feel you do not apply to this law. The cops of our cities are doing their job by stopping people who does not follow this law. The main reason why this show not be condoned is, because most of these situations a citizen could get in a car accident. Approximately 1.6 million citizens die due to being on their cell phone while operating a vehicle. Nearly 390,000 people get real serious injuries due to texting and driving. One out of four of every car accident in the United States are due to texting and driving. In my opinion I believe texting and driving is the main reason we have so many deaths every year. texting and driving needs to be/ stay illegal. That way people will not do it so often, which could decrease the death rate each year. Texting and driving could possibly be the reason why our death rates increase over time. Especially in this day and age, because almost everything now is technology based I believe that texting and driving is not safe for yourself or others, so technology should be the last thing you are on while behind and active vehicle. Texts, calls, emails all can wait until you are in a secure place (not behind the wheel) to respond, and also that way nobody could die or have an injury. |
| Prompt |
Consider the following problem statement: Explain the reasons why distractions while driving, specifically texting and driving, should be illegal and the consequences of not following this law. Use statistics and personal opinion to support your argument. Write a well-structured essay of approximately 325 words expressing your clear opinion on the issue. Your essay should provide a thorough analysis of the problem, explore potential implications, and propose viable solutions or recommendations. Be sure to consider different perspectives surrounding the problem and support your opinions with relevant evidence, examples, or logical reasoning. Additionally, emphasize the importance of addressing the issue and the potential impact of implementing your proposed solutions. Essay: |
| Machine |
Distractions while driving are a multi-faceted danger, with texting and driving sitting at the proverbial epicenter. The amalgamation of visual, manual, and cognitive distractions involved in this act makes it a leading hazard on the roads. This essay elucidates the imperative nature of legislation against texting and driving, buttressed by statistics and personal outlook, and explores the ramifications of non-compliance alongside potential resolutions. To initiate, consider the statistical landscape painting the grim reality. The National Highway Traffic Safety Administration (NHTSA) reports that in the United States alone, approximately 3,000 people lose their lives annually in crashes attributed to distracted driving. Texting is particularly pernicious; it takes the driver2̆019s eyes off the road for an average of five seconds, amounting to blind traversal of the length of a football field at 55 miles per hour. This nexus of inattention and control relinquishment is the breeding ground for tragedies. Moreover, the moral gravity of the situation cannot be overstated. When a driver engages in texting, they tacitly place their transient communication above the safety of fellow road users. This reflects a negligent societal value towards human life which by any humane standard is inexcusable. Hence, illegalizing such behavior manifests a collective ethos valuing life and responsibility. However, legal proscription must be complimented by earnest enforcement. Weak adherence to anti-texting laws, muddled by insubstantial sanctions, can subvert the intended deterrent effect. I assert that the consequence for distracted driving should encompass hefty fines, driving license suspensions, and, for repeat offenders or resulting in severe injury or death, significant custodial sentences. To mitigate the urge to reach for the phone, technological solutions like smartphone applications that disable texting while driving, and robust in-vehicle systems, need broader promotion. Educational campaigns fortifying public awareness about the consequences of distracted driving, seem pivotal to shifting attitudes, especially among impressionable young drivers. A symbiosis of legal strictures, technology, and education presents a tripartite solution to this problem. In conclusion, driving demands undivided attention, and texting fractures this focus. Legislating against such distractions is not only a proactive step towards reducing traffic-related fatalities and injuries, but it also reinforces a societal norm that prioritizes collective safety over momentary convenience. The significance of addressing this issue is underscored by the unbearable cost of human lives that could have been saved. In the grand tableau of road safety, cutting down the text during driving is not just a law; it is an ethical imperative that should carry the weight of universal endorsement. |
| Language | English |
| LLM Generator | GPT-4 |
| Domain | arXiv |
| Title |
The Spitzer c2d Survey of Large, Nearby, Insterstellar Clouds. IX. The Serpens YSO Population As Observed With IRAC and MIPS |
| Human |
We discuss the results from the combined IRAC and MIPS c2d Spitzer Legacy observations of the Serpens star-forming region. In particular we present a set of criteria for isolating bona fide young stellar objects, YSO’s, from the extensive background contamination by extra-galactic objects. We then discuss the properties of the resulting high confidence set of YSO’s. We find 235 such objects in the 0.85 deg2̂ field that was covered with both IRAC and MIPS. An additional set of 51 lower confidence YSO’s outside this area is identified from the MIPS data combined with 2MASS photometry. We describe two sets of results, color-color diagrams to compare our observed source properties with those of theoretical models for star/disk/envelope systems and our own modeling of the subset of our objects that appear to be star+disks. These objects exhibit a very wide range of disk properties, from many that can be fit with actively accreting disks to some with both passive disks and even possibly debris disks. We find that the luminosity function of YSO’s in Serpens extends down to at least a few x .001 Lsun or lower for an assumed distance of 260 pc. The lower limit may be set by our inability to distinguish YSO’s from extra-galactic sources more than by the lack of YSO’s at very low luminosities. A spatial clustering analysis shows that the nominally less-evolved YSO’s are more highly clustered than the later stages and that the background extra-galactic population can be fit by the same two-point correlation function as seen in other extra-galactic studies. We also present a table of matches between several previous infrared and X-ray studies of the Serpens YSO population and our Spitzer data set. |
| Prompt |
Write an abstract of an article from ArXiv given its title so that it is indistinguishable from a genuine human-written abstract. Write in the style that most abstracts from ArXiv are written. Title: The Spitzer c2d Survey of Large, Nearby, Insterstellar Clouds. IX. The Serpens YSO Population As Observed With IRAC and MIPS. Abstract: |
| Machine |
This paper presents the results of the Spitzer c2d survey of the Serpens star-forming region, focusing on the young stellar object (YSO) population as observed with the Infrared Array Camera (IRAC) and the Multiband Imaging Photometer for Spitzer (MIPS). The survey identified a significant number of YSO candidates, providing a comprehensive view of star formation in this region. The data obtained from IRAC and MIPS observations were used to construct color-color and color-magnitude diagrams, which were then utilized to classify the YSOs into different evolutionary stages. The spatial distribution of these objects was also analyzed, revealing a correlation with the dense molecular gas in the region. The results contribute to our understanding of the star formation process in large, nearby, interstellar clouds and provide a valuable resource for future studies. |
| Language | Arabic |
| LLM Generator | ChatGPT |
| Domain | News |
| Title |
{RLtext} الجيش التركي يعلن جاهزيته للعبور إلى ليبيا.. المعارك تتجدد بطرابلس وقوات حفتر تتراجع |
| Human |
{RLtext} تجددت المعارك اليوم الجمعة بمحيط العاصمة الليبية طرابلس، وأفادت التقارير |
| بتراجع قوات اللواء المتقاعد خليفة حفتر ومقتل أحد أبرز قادتها، في حين أكد الجيش | |
| التركي جاهزيته للتوجه لليبيا فور تلقيه التعليمات. وقد أفاد مصدر عسكري من حكومة | |
| الوفاق بتجدد الاشتباكات المسلحة بأغلب محاور القتال جنوبي طرابلس. واستعادت قوات | |
| الوفاق مواقع خسرتها في وقت سابق، وأجبرت قوات حفتر على الانسحاب إلى خلف معسكر | |
| اليرموك بمحيط طرابلس. وذكرت مصادر من قوات الوفاق أن مقاتلي الخطوط الأمامية في | |
| قوات حفتر كانوا من مسلحي شركة ’’فاغنر‘‘ الروسية، الذين ينفذون عمليات قتالية ضد | |
| قوات حكومة الوفاق باستخدام أسلحة نوعية حديثة. وأكدت وسائل إعلام موالية لقوات | |
| حفتر أنها فقدت قائدا عسكريا بارزا يحمل رتبة عقيد ركن ويدعى فتحي محمد المسلاتي. | |
{RLtext}ولقي المسلاتي مصرعه اليوم الجمعة بقصف صاروخي في محور اليرموك جنوب العاصمة |
| طرابلس. وكان المسلاتي يقود سلاح الدبابات والمدرعات باللواء 106 التابع لقوات | |
| حفتر. في السياق ذاته، استهدفت طائرات حفتر اليوم الجمعة تمركزات عسكرية بمحيط | |
| مصفاة الزاوية شمال غربي البلاد. وكانت المدينة قد تعرضت أمس الخميس لثلاث غارات | |
| جوية أسفرت إحداها عن مقتل مدنيين اثنين وجرح ثمانية آخرين نتيجة سقوط صواريخ على | |
| محلات عدة بالطريق الساحلي في الزاوية. واليوم الجمعة خرجت في المدينة مسيرة تندد | |
| بغارات قوات حفتر التي تستهدف المنشآت النفطية وتؤدي إلى سقوط ضحايا من المدنيين. | |
| وعبر المتظاهرون عن رفضهم للهجوم الذي تشنه قوات حفتر على العاصمة طرابلس، ونددوا | |
| بمواقف الدول الداعمة له. في انتظار الأوامرمن جانبها، أكدت مصادر عسكرية في | |
| أنقرة أن القوات المسلحة التركية مستعدة للتوجه إلى ليبيا والقيام بمهامها حال | |
| تلقيها التعليمات. جاء ذلك عقب مؤتمر صحفي للمتحدثة باسم وزارة الدفاع ناديدا | |
| شبنم أكتوب قدمت فيه تقييما لأنشطة الوزارة خلال العام 2019. وعن التوجه إلى | |
| ليبيا، قالت المصادر إن القوات المسلحة التركية مستعدة للقيام بكافة أنواع المهام | |
| داخل وخارج تركيا حال تلقيها التعليمات. وأمس الخميس، توقع الرئيس التركي رجب طيب | |
| أردوغان حصول حكومته على تفويض من البرلمان في الثامن من يناير/كانون الثاني | |
| المقبل من أجل إرسال جنود إلى ليبيا، تلبية لدعوة الحكومة الشرعية. وأكد أن بلاده | |
| ستقدم جميع أنواع الدعم لحكومة طرابلس في كفاحها ضد "الجنرال الانقلابي خليفة | |
| حفتر، المدعوم من دول أوروبية وعربية مختلفة". وفجر الخميس دخلت حيز التنفيذ | |
| مذكرة التفاهم الخاصة بالتعاون الأمني والعسكري بين تركيا وليبيا بعد نشرها في | |
| الجريدة الرسمية التركية. وفي وقت سابق، قال السفير الليبي في تركيا عبد الرزاق | |
| مختار عبد القادر إن بلاده تحتاج دعما عسكريا في الدفاع الجوي والتدريب الخاص، | |
| وذلك في ظل أنباء عن تقديم حكومة الوفاق طلبا رسميا إلى أنقرة للحصول على دعم | |
| عسكري بحري وبري وجوي. ومنذ 4 أبريل/نيسان الماضي، يشهد محيط طرابلس معارك مسلحة | |
| بين قوات حكومة الوفاق المعترف بها دوليا وقوات الجنرال المتقاعد خليفة حفتر. | |
| Prompt |
[…]
| Make sure the content is around 385 words. | |
| Title: |
[…]
Language Arabic LLM Generator Jais Domain News Title {RLtext} الجيش التركي يعلن جاهزيته للعبور إلى ليبيا.. المعارك تتجدد بطرابلس وقوات حفتر تتراجع Human {RLtext} تجددت المعارك اليوم الجمعة بمحيط العاصمة الليبية طرابلس، وأفادت التقارير بتراجع قوات اللواء المتقاعد خليفة حفتر ومقتل أحد أبرز قادتها، في حين أكد الجيش التركي جاهزيته للتوجه لليبيا فور تلقيه التعليمات. وقد أفاد مصدر عسكري من حكومة الوفاق بتجدد الاشتباكات المسلحة بأغلب محاور القتال جنوبي طرابلس. واستعادت قوات الوفاق مواقع خسرتها في وقت سابق، وأجبرت قوات حفتر على الانسحاب إلى خلف معسكر اليرموك بمحيط طرابلس. وذكرت مصادر من قوات الوفاق أن مقاتلي الخطوط الأمامية في قوات حفتر كانوا من مسلحي شركة "فاغنر" الروسية، الذين ينفذون عمليات قتالية ضد قوات حكومة الوفاق باستخدام أسلحة نوعية حديثة. وأكدت وسائل إعلام موالية لقوات حفتر أنها فقدت قائدا عسكريا بارزا يحمل رتبة عقيد ركن ويدعى فتحي محمد المسلاتي. ولقي المسلاتي مصرعه اليوم الجمعة بقصف صاروخي في محور اليرموك جنوب العاصمة طرابلس. وكان المسلاتي يقود سلاح الدبابات والمدرعات باللواء 106 التابع لقوات حفتر. في السياق ذاته، استهدفت طائرات حفتر اليوم الجمعة تمركزات عسكرية بمحيط مصفاة الزاوية شمال غربي البلاد. وكانت المدينة قد تعرضت أمس الخميس لثلاث غارات جوية أسفرت إحداها عن مقتل مدنيين اثنين وجرح ثمانية آخرين نتيجة سقوط صواريخ على محلات عدة بالطريق الساحلي في الزاوية. واليوم الجمعة خرجت في المدينة مسيرة تندد بغارات قوات حفتر التي تستهدف المنشآت النفطية وتؤدي إلى سقوط ضحايا من المدنيين. وعبر المتظاهرون عن رفضهم للهجوم الذي تشنه قوات حفتر على العاصمة طرابلس، ونددوا بمواقف الدول الداعمة له. في انتظار الأوامرمن جانبها، أكدت مصادر عسكرية في أنقرة أن القوات المسلحة التركية مستعدة للتوجه إلى ليبيا والقيام بمهامها حال تلقيها التعليمات. جاء ذلك عقب مؤتمر صحفي للمتحدثة باسم وزارة الدفاع ناديدا شبنم أكتوب قدمت فيه تقييما لأنشطة الوزارة خلال العام 2019. وعن التوجه إلى ليبيا، قالت المصادر إن القوات المسلحة التركية مستعدة للقيام بكافة أنواع المهام داخل وخارج تركيا حال تلقيها التعليمات. وأمس الخميس، توقع الرئيس التركي رجب طيب أردوغان حصول حكومته على تفويض من البرلمان في الثامن من يناير/كانون الثاني المقبل من أجل إرسال جنود إلى ليبيا، تلبية لدعوة الحكومة الشرعية. وأكد أن بلاده ستقدم جميع أنواع الدعم لحكومة طرابلس في كفاحها ضد "الجنرال الانقلابي خليفة حفتر، المدعوم من دول أوروبية وعربية مختلفة". وفجر الخميس دخلت حيز التنفيذ مذكرة التفاهم الخاصة بالتعاون الأمني والعسكري بين تركيا وليبيا بعد نشرها في الجريدة الرسمية التركية. وفي وقت سابق، قال السفير الليبي في تركيا عبد الرزاق مختار عبد القادر إن بلاده تحتاج دعما عسكريا في الدفاع الجوي والتدريب الخاص، وذلك في ظل أنباء عن تقديم حكومة الوفاق طلبا رسميا إلى أنقرة للحصول على دعم عسكري بحري وبري وجوي. ومنذ 4 أبريل/نيسان الماضي، يشهد محيط طرابلس معارك مسلحة بين قوات حكومة الوفاق المعترف بها دوليا وقوات الجنرال المتقاعد خليفة حفتر.
Prompt {RLtext} بناءً على التفاصيل التالية لمقالة أخبار عربية، قم بإنشاء المحتوى باللغة العربية .تأكد من أن المحتوى يتراوح حول 385 كلمات. العنوان: الجيش التركي يعلن جاهزيته للعبور إلى ليبيا.. المعارك تتجدد بطرابلس وقوات حفتر تتراجع الوصف: تجددت المعارك اليوم الجمعة بمحيط العاصمة الليبية طرابلس، وأفادت التقارير بتراجع قوات اللواء المتقاعد خليفة حفتر ومقتل أحد أبرز قادتها، بينما أكد الجيش التركي جاهزيته للتوجه لليبيا فور تلقيه التعليمات. المرجع: المصدر : الجزيرة + وكالات المحتوى: Machine {RLtext} الجيش التركي يعلن جاهزيته للعبور إلى ليبيا.. المعارك تتجدد بطرابلس وقوات حفتر تتراجع
أفادت التقارير الواردة من محيط العاصمة الليبية طرابلس بتجدد الاشتباكات اليوم الجمعة بين قوات حكومة الوفاق الوطني المعترف بها دوليا وقوات اللواء المتقاعد خليفة حفتر. كما أفادت التقارير بتراجع قوات حفتر ومقتل أحد أبرز قادتها.
وفي تطور آخر، أعلن الجيش التركي عن جاهزيته للعبور إلى ليبيا ونشر قواته فور تلقيه التعليمات من الحكومة التركية. وجاء هذا الإعلان بعد اجتماع للأمن القومي التركي يوم الخميس، برئاسة الرئيس رجب طيب أردوغان.
وكانت قوات حفتر قد حاولت في وقت سابق السيطرة على طرابلس، لكنها واجهت مقاومة شديدة من القوات الموالية للحكومة المعترف بها دوليا. وأسفرت المعارك الدائرة منذ أبريل / نيسان الماضي عن مقتل المئات وتشريد الآلاف.
من جهة أخرى، أكد الرئيس التركي رجب طيب أردوغان أن بلاده لن تظل صامتة إزاء قوات حفتر التي تدعمها الإمارات ومصر وروسيا. واتهم أردوغان هذه الدول بالتورط في سفك الدماء في ليبيا.
وفي سياق متصل، قال رئيس الحكومة الإيطالية جوزيبي كونتي إن بلاده تشعر بالقلق إزاء التطورات في ليبيا، مؤكدا أن ليبيا تحتاج إلى فعل وليس أقوال.
وتحاول الأمم المتحدة التوسط بين الأطراف المتناحرة في ليبيا لإنهاء القتال الدائر في البلاد. لكن المنظمة الدولية لم تتمكن من التوصل إلى تسوية سياسية حتى الآن.
وقد حذرت منظمات دولية من بينها الأمم المتحدة ومنظمة العفو الدولية من أن القتال في ليبيا قد وصل إلى مستويات جديدة من الرعب مع استهداف المدنيين والمدارس والمستشفيات.
ولا تزال الأوضاع في ليبيا تشهد تصعيدا على الأرض رغم الجهود الدولية لوقف إطلاق النار. ولا تزال قوات حفتر مستمرة في حصارها لطرابلس للشهر السادس على التوالي.
المراجع: - المصدر : الجزيرة + وكالات
Language German LLM Generator ChatGPT Domain Wikipedia Title Abwasserpumpe Human Schmutzwasserpumpen sind in der Lage, auch grob verunreinigtes Wasser (aber auch andere fluide Medien) mit hohem Feststoffanteil wie Steinen, Geröll und Ähnlichem zu fördern. Sie arbeiten in der Regel nicht selbstansaugend, sondern sind als Tauchpumpe ausgelegt, bei denen ein Schaufelrad das Fluid fördert. Dafür werden besondere Laufradformen wie Freistromrad, Einkanalrad, Mehrkanalrad, Diagonalrad, Schraubenrad wie auch Propellerrad verwendet. Kleinere Pumpen haben meistens ein Freistromrad, da dies am wenigsten störanfällig ist und einen großen freien Durchgang für die Feststoffe bietet. Pumpen mit Freistromrädern haben jedoch schlechte Wirkungsgrade und erreichen auch keine großen Förderhöhen; deshalb wählt man bei größeren Maschinen meistens andere Laufradformen. Schmutzwasserpumpen sind auf die Fördermenge ausgelegt – typische Förderleistungen von 5.000 bis 15.000 Liter pro Stunde (etwa 3 bis 5 Liter je Sekunde) – und weniger auf den Druck, die Förderhöhe beträgt häufig nur einige Meter. Der Antrieb erfolgt in der Regel durch einen Elektromotor, je nach Anwendungsfall aber auch mittels eines Verbrennungsmotors erfolgen. Eingesetzt wird dieser Pumpentyp etwa im Tiefbau zum Auspumpen von Baugruben, im Hochwasserschutz und in der Katastrophenhilfe, zur Bewässerung in Landwirtschaft und Gartenbau oder in der Landschaftsarchitektur zur Befüllung von künstlichen Wasseranlagen. Kleinere Modelle werden auch im Heimbereich verwendet, um das Wasser für den Rasensprenger nicht dem „teuren“ Trinkwasser zu entnehmen, sondern aus nahen Fließgewässern oder Teichen (hierbei sind aber wasserrechtliche Vorschriften zu beachten). Prompt This is a sample German Wikipedia summary section for the title "WBDJ": Der sozialistisch ausgerichtete Weltbund der Demokratischen Jugend (WBDJ) wurde am 10. November 1945 auf der Weltjugendkonferenz in London gegründet. Er ist von den Vereinten Nationen als internationale nichtstaatliche Jugendorganisation anerkannt.
Dem WBDJ gehören über 150 Mitgliedsorganisationen (nach eigenen Angaben 2012 über 270) aus 109 Ländern an, die rund 30 Millionen Jugendliche vertreten. Diese Jugendorganisationen sind kommunistisch oder sozialistisch orientiert und/oder sie gehören Befreiungsbewegungen an.
Der WBDJ ist der Hauptträger der in unregelmäßigen Abständen seit 1947 stattfindenden Weltfestspiele der Jugend und Studenten.Der WBDJ war eine kommunistische Frontorganisation. In Folge des Zusammenbruchs des sozialistischen Lagers geriet der WBDJ 1990/91 in eine schwere Krise. Der künftige Charakter der Vereinigung wurde in Frage gestellt. Der WBDJ überstand jedoch diese Krise und stellte seine reguläre Arbeitsweise wieder her. 1997 wurden erstmals wieder Weltfestspiele auf Kuba durchgeführt. Höchstes Organ des WBDJ ist die Generalversammlung (bis 1957 Weltkongress), die zuletzt 2019 in Nikosia stattfand. Diese war mit 160 Vertretern von 93 Organisationen die größte Generalversammlung der letzten 30 Jahre.
Using a similar formatting and style, write a German Wikipedia summary section for the title "Abwasserpumpe". The summary should have at least 250 words. Machine Abwasserpumpe ist eine elektrische Pumpe zur Förderung von Abwasser und Schmutzwasser. Zusätzlich muss sie in der Lage sein, Feststoffe wie Haare, Papier oder Faserstoffe zu zerkleinern, um Verstopfung zu vermeiden. Man zählt die Abwasserpumpe zu den "Tauchmotorpumpen", da das gesamte Gerät unter Wasser getaucht wird. Das Pumpwerk (bestehend aus Saugkorb, Hydroscheider und Laufrad) wird dabei in einer Pumpstation eingesetzt. Der Antrieb des Laufrads geschieht durch einen Elektromotor. Die Pumpe trägt dazu bei, dass Abwasser und Schmutzwasser von der Einleitungsstelle zu den Kläranlagen oder sammelnden Kanalsystemen weitergeleitet werden. Insbesondere in Privathaushalten und Großküchen ist die Abwasserpumpe nützlich, da sie den Transport von Abwasser wesentlich erleichtert. Allerdings sollte man auf Handelsnamen achten: nicht alle Tauchmotorpumpen sind auch für den Einsatz in der Abwasserverarbeitung zugelassen.
Eine häufig auftretende Schwierigkeit bei der Abwasserpumpe ist der sogenannte "Longpumpen". Hierbei kann der Abwasserstrom nicht mehr vollständig durch den Pumpenkopf geleitet werden, wodurch sich der Anteil sedimenter Stoffe erhöht. Darüber hinaus kann lang anhaltender Stress auf wenige Teile der Pumpe - etwa die Achse oder das Laufrad - schnell zu Abnutzungserscheinungen führen. Regelmäßige Wartung und Einhaltung von Grenzwerten für die Höhe des Abwasserdrucks oder die Größe der durchgeleiteten Partikel können diese Probleme vermindern.
Die Qualität von Abwasserpumpen lässt sich anhand des maximalen zulässigen Drucks, der maximale Durchflussmenge, dem verwendeten Material oder auch in Hinblick auf Effizienz merklich voneinander unterscheiden. Es kann auch sinnvoll sein, beim Kauf einer Abwasserpumpe auf vorhandene Siegel und Zertifikate wie CE, VDE oder TÜV zu achten.
Language German LLM Generator ChatGPT Domain News Title Journalisten und Publikum sägen an den Ästen, auf denen sie sitzen Human Verlogen, unglaubwürdig, verwöhnt: So werden Journalisten heute taxiert. Sie werden alle in einen Topf geworfen, ihre Arbeit scheint wertlos, beliebig und ein Ärgernis. Das muss aufhören. Medienpublikum, Medienmanager, Medienforscher und Journalisten sägen derzeit selber an den Ästen, auf denen sie sitzen. Gewiss gibt es Journalismus, der über die Stränge schlägt oder abgehoben ist. Und es gibt zu viele Journalisten, die selbstgefällig sind, keine Fehler zugeben und sich diebisch freuen, wenn sie einen anderen in die Pfanne hauen können. Das sei nicht kleingeredet, aber ins Verhältnis gesetzt. Wir wissen, dass es Banker gibt, die in die eigene Tasche wirtschaften und zahllose Kunden ruinieren, Automobilclubs, die Ranglisten manipulierten, Unternehmer, die mit illegalen Abschalteinrichtungen Abgasnormen umgingen, Politiker, die politische Gegner ausspionierten, korrupte Fußballfunktionäre, quacksalbernde Ärzte, Winkeladvokaten. Und wir wissen das alles oft gerade, weil es eben neben schwarzen Schafen noch andere Journalisten gibt. Ende Februar wurde Spotlight mit einem Oscar ausgezeichnet: ein Film über eine Recherche, die die Missbrauchsfälle der katholischen Kirche in Boston enthüllte. Für diese Recherche gewann das Investigativteam des Boston Globe Jahre zuvor einen Pulitzerpreis. Doch nicht nur der Enthüllung von Missständen wegen müssen wir eine Lanze brechen für einen aufklärenden, verantwortungs- und werteorientierten Informationsjournalismus. Wir brauchen Medien, die uns Arenen öffnen, in denen wir um die besten Argumente ringen und diskutieren, wie wir künftig in unserer Gesellschaft leben wollen. Hören wir auf, an Ästen zu sägen, und fordern wir einen Journalismus ein, der sich bewusst seinem Publikum zuwendet, der Zweifel, Ängste und Sorgen ernst nimmt und auch konstruktiv aufzeigt, wie Probleme gelöst werden können. Prompt Schreibe einen deutschen Nachrichtenartikel zu einem vorgegebenen Titel. Der Artikel muss mindestens 250 Wörter und höchstens 500 Wörter enthalten. Schreibe ausschließlich den Artikel und keine Anweisungen, Kommentare, Werbeanzeigen, Hashtags oder ähnliches.
Titel: Journalisten und Publikum sägen an den Ästen, auf denen sie sitzen
Artikel: Machine Heute leben wir in der Welt des Internets. Dies hat dazu beigetragen, dass jeder seinen Standpunkt darlegen und seine Meinung veröffentlichen kann. Aber haben wir dadurch eine Situation geschaffen, die zur Unabhängigkeit führt, oder haben wir uns in ein Gefängnis verwandelt, das von einer Gruppe von Menschen kontrolliert wird, die unsere Freiheit zensieren und unsere Perspektive beeinflussen?
In den Tageszeitungen sind Journalisten in einer ähnlichen Situation gefangen. In dem Bestreben, die Meinungsfreiheit zu fördern und Nachrichten zu verbreiten, wird auch diese Gruppe von Menschen von der Unabhängigkeit des Internets gefangen genommen. Aber es gibt eine große Lücke zwischen der Zensur im Internet und der Zensur bei Tageszeitungen.
Es ist wahr, dass Tageszeitungen dazu neigen, sich auf die Meinungen ihrer Leser zu konzentrieren, um einen zufriedenstellenden Inhalt bereitzustellen, der alle belohnt. Aber es ist auch wahr, dass diese Eigenarten der Tageszeitungen, als informelles Werkzeug für die politische Agenda genutzt werden kann. Werbewirtschaftliche und politisch motivierte Entscheidungen zwingen Journalisten dazu, die Wahrheit von dem zu unterscheiden, was derzeit trendig ist.
Die Leserschaft selbst, die mit Zeitungen und anderen Publikationen sozialisiert wurde, erwartet manchmal einen sehr spezifischen Nachrichteninhalt, der möglicherweise mehr mit ihrer eigenen ideologischen Perspektive als mit einer objektiven Wahrheit übereinstimmt. Eine solche Publikumsbindung kann dazu führen, dass Journalisten gezwungen werden, in eine bestimmte Richtung zu tendieren und sich zwangsläufig von der Wahrheit zu entfernen.
Schließlich werden Journalisten dazu gezwungen, Artikel zu schreiben, die das Publikum möglicherweise nicht hören will. Sie schreiben über schwierige Themen wie Armut, Arbeitslosigkeit, Drogenmissbrauch, Hunger und Kriege. Aber einige Journalisten sind sich nicht sicher, wie sie über diese Probleme schreiben sollen, ohne das Publikum nicht zu verlieren, das dazu tendiert, einfache Geschichten mit einer positiveren Note zu bevorzugen. Dies kann dazu führen, dass Journalisten eine eingeschränkte Rolle in der Informationsbeschaffung übernehmen und das Publikum den verbleibenden Raum besetzt.
Das Problem der Journalisten und Publikum, die an den Ästen sägen, auf denen sie sitzen, ist also ein beunruhigender Trend, der unsere Freiheit und Möglichkeiten beeinträchtigen kann. Wenn wir befreit und unabhängig bleiben wollen, benötigen wir mehr professionelle und mutige Journalisten, die bereit sind, eine Vielzahl von Lösungen für unser komplexes Leben zu präsentieren, unabhängig von den eigenen Interessen. Die Herausforderung besteht darin, das kritische Denken und die Unabhängigkeit der Medien zu fördern, um sicherzustellen, dass unsere Freiheit und unsere Weltsicht nicht von einer einflussreichen Gruppe von Menschen dominiert werden. Wir sollten das Bestreben der Journalisten in ihrem Bemühen um Unabhängigkeit unterstützen und uns dafür einsetzen, damit jeder die Freiheit hat, seine Meinung zu äußern und die Wahrheit zu suchen.
Language Italian LLM Generator camoscio-70B Domain News Human Francesco Schettino rilancerà il turismo delle Marche? Non c’è ancora nessuna ufficialità, ma già è scoppiata la bufera. Infatti, léx comandante della Costa Concordia, secondo alcuni quotidiano locali, sarebbe pronto a diventare testimonial di Numana per il rilancio del turismo nella Riviera del Conero. Il sindaco però si è subito affrettato a precisare che si tratta di una idea malsana e fantasiosa. Sarà. Intanto però il legale di Schettino ha annunciato che il suo cliente avrà un colloquio di lavoro nei prossimi giorni. " Auspichiamo vivamente che questa buia fantasia non abbia fondamento alcuno: una persona divenuta famosa per un evento così negativo sicuramente non deve essere associata a una località che vive di turismo, bella e solare ", ha tuonato Carlo Neumann , presidente dell’Associazione Riviera del Conero. "Apprezzo lo sforzo di fantasia, ma bisogna partorire qualcosa di più interessante ", ha commentato il sindaco di Numana, Marzio Carletti (Pd). Insomma, ancora non si sa se Schettino metterà piede a Numana. Quello che è indubbio è che non ci sarà nessun inchino ad attenderlo Prompt ### Istruzione: Dato il testo ’Schettino testimonial per rilanciare il turismo nella Riviera del Conero?’ scrivete un articolo di almeno 1000 parole in Italiano di cui quello è il titolo. ### Risposta: Machine "La Riviera del Conero, situata sulla costa adriatica della regione Marche dell’Italia centrale, è una delle destinazioni turistiche più popolari d’Europa. È conosciuta per le sue spiagge di sabbia dorata e acque cristalline, così come per i suoi pittoreschi villaggi mediterranei che si snodano lungo la riva. Tuttavia, negli ultimi anni il numero di visitatori nella zona è diminuito significativamente a causa della crisi economica globale. Per cercare di invertire questa tendenza negativa, l’amministrazione locale ha deciso di rivolgersi al capitano Francesco Schettino, ex comandante della nave da crociera Costa Concordia naufragata nel 2012, per aiutare a promuovere nuovamente la Riviera del Conero come meta vacanziera ideale. Secondo gli amministratori locali, poiché il capitano Schettino è diventato un volto familiare in tutto il mondo grazie alla sua associazione con la tragedia della Costa Concordia, potrebbe essere uno strumento efficace per attirare attenzione sulla bellezza naturale ed esotica della Riviera del Conero"